At the end of a race a runner decelerates from a velocity of 11.00 m/s at a rate of 0.300 m/s2.

How far does she travel in the next 15.0 s? =m

What is her final velocity? =m/s
1
Expert's answer
2018-09-10T12:51:09-0400
The distance
S=v_i t-(at^2)/2=11.000×15.0-(0.300×〖15.0〗^2)/2=131.25 m
The final velocity

The final velocity
v_f=v_i-at=11.000-0.300×15.0=6.50 m/s